63.5 mol of p4o10 contains how many moles of P
Answer #1
Concepts and reason

Mole concept:  Mole is the amount of substance of a system, which contains atoms and molecules. One mole of any substance or element contains exactly 6.023x102
molecules or atoms of that substance. This number of units is called Avogadro’s number. According to Avogadro’s hypothesis, 1 mole
 of the substance occupies 22.4L
 at standard temperature and pressure.

Number of atoms = Avogadro’s number (6.023×1023){\rm{(6}}{\rm{.023 \times 1}}{{\rm{0}}^{{\rm{23}}}}{\rm{)}} .

Avogadro’s number defines the number of atoms present in one mole of a substance.

Therefore, 1mole = 6.023×1023{\rm{6}}{\rm{.023 \times 1}}{{\rm{0}}^{{\rm{23}}}} atoms.

Atoms are made up of tiny particles in which a positive charged particle in an atom is called proton, negative charged particle is called electron, and the neutral charged (No charge) particle is called neutrons.

Fundamentals

By using mass and molar mass of substance, moles can be calculated.

Moles=massmolecularmass{\rm{Moles}}\,\,{\rm{ = }}\,\,\,\frac{{{\rm{mass}}}}{{{\rm{molecular}}\,\,{\rm{mass}}}}

The total number of atoms present in a substance can be calculated by multiplying the moles with Avogadro’s number (6.023×1023){\rm{(6}}{\rm{.023 \times 1}}{{\rm{0}}^{{\rm{23}}}}{\rm{)}} .

The compound P4O10{{\rm{P}}_{\rm{4}}}{{\rm{O}}_{{\rm{10}}}} contains ten moles of oxygen and four moles of phosphorous. Thus, the number of moles of phosphorous in one mole of P4O10{{\rm{P}}_{\rm{4}}}{{\rm{O}}_{{\rm{10}}}} is four.

The number of moles of phosphorous in 63.5mol{\rm{63}}{\rm{.5 mol}} of P4O10{{\rm{P}}_{\rm{4}}}{{\rm{O}}_{{\rm{10}}}} is identified as follows:

Numberofmolesin63.5mol=63.5×numberofmolesinonemol=63.5×4=254moles\begin{array}{c}\\{\rm{Number of moles in 63}}{\rm{.5 mol}}\,{\rm{ = }}\,63.5{\rm{ \times number of moles in one mol}}\\\\{\rm{ = }}\,63.{\rm{5 \times 4}}\\\\{\rm{ = }}\,254{\rm{ moles}}\\\end{array}

Ans:

The number of moles of phosphorous in 63.5mol{\bf{63}}{\bf{.5 mol}} of P4O10{{\bf{P}}_{\bf{4}}}{{\bf{O}}_{{\bf{10}}}} is 254moles{\bf{254 moles}} .